Find lim x tends to 1 [x^x-1/xlogx]

Grade:11

1 Answers

Abhishek Doyal
10 years ago
when lim tends to 1 it's become 0/0 form

[x^x-1/xlogx] differentiation of x^x-1=x^x(1+log(x))

differentiation ofxlogx=1+log (x)


so by L hospital rule differentiate the fxn
y=x^x(1+log(x))/1+log(x)
put x=1 y=1
limx tends to1=1 ans


Abhishek doyal
IIT Bombay
Askiitian tutor